3x - 2y = 8
-12 + 8y = 32

Answer by Alan3354(69443)

3x - 2y = 8
-12 + 8y = 32 (the x is missing)
--------------
3x - 2y = 8 eqn 1
-12x + 8y = 32 eqm 2
Divide eqn 2 by -4
3x - 2y = -8
There's no solution, 3x-2y cannot = both 8 and -8